OpenStudy (anonymous):

Solve the Proportion . 15t 2t + 3 ------- = ------------ 5 6 A. 0.03 B. 0.1875 C. 0.0375 D. 0.15

OpenStudy (kash_thesmartguy):

Welcome to OpenStudy! Here, you can enjoy by learning. Here is the legendary tutorial for our legendary OpenStudiers: http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8 Created by: @Kash_TheSmartGuy

OpenStudy (campbell_st):

if the question is \[\frac{15t}{5} = \frac{2t + 3}{6}\] cross multiply 90t = 10t + 15 then solve for t

OpenStudy (anonymous):

I got 7.5t which isn't a choice

OpenStudy (campbell_st):

wow.... I can understand why 7.5t isn't a choice... 90t = 10t + 15 subtract 10t from both sides of the equation -10t -10t ------------------- 80t = 15 now solve for t

OpenStudy (anonymous):

Well I was taught to do the constant first which would be subtracting 15 from each side

OpenStudy (anonymous):

And then divide the coefficient